# Quantum determinism and completeness restored by indistinguishability and long-time particle detection

### v1

@@ Abstract
We argue that measurement data in quantum physics can be rigorously interpreted only as a result of a statistical, macroscopic process, taking into account the indistinguishable character of identical particles. Quantum determinism is in principle possible on the condition that a fully-fledged quantum model is used to describe the measurement device in interaction with the studied object as one system. In contrast, any approach that relies on Born’s rule discriminates the dynamics of a quantum system from that of the detector with which it interacts during measurement. In this work, we critically analyze the validity of this measurement postulate applied to single-event signals. In fact, the concept of “individual” particle becomes inadequate once both indistinguishability and a scattering approach allowing an unlimited interaction time for an effective detection, are considered as they should be, hence preventing the separability of two successive measurement events. In this context, measurement data should therefore be understood only as a result of statistics over many events. Accounting for the intrinsic noise of the sources and the detectors, we also show with the illustrative cases of the Schrödinger cat and the Bell experiment that once the Born rule is abandoned on the level of a single particle, realism, locality and causality are restored. We conclude that indiscernibility and long-time detection process make quantum physics not fundamentally probabilistic.
@@ I Introduction
Here, we argue that the Born rule (or probability measurement postulate) is a non-essential mathematical convenience that should no longer be used, even though for nearly 100 years, much research in quantum physics has relied upon it for the interpretation of quantum experimental data. In a nutshell, this rule states that the probability density of finding a system in a given state, immediately after measurement, is proportional to the square of the amplitude of the system’s wave function in that state immediately before measurement, thus assuming that the system was isolated from its environment during the measurement process. In this framework, the measurement of a physical quantity can only yield one of the eigenvalues of the corresponding observable with a certain probability. As such, the Born rule separates the dynamics of the studied quantum system from that of the measurement device, thus assuming that measurement data correspond to observables as properties of the studied quantum system only. However, the elephant in the room that has been systematically overlooked in practice, is that there is no reason why quantum mechanics should discriminate between the object of study and the measurement device it is in interaction with as though they were independent entities. A precise knowledge of the interaction between the detector and the quantum system is required for a detailed and rigorous description of the measurement process and hence a sound interpretation of the data. This is a considerable challenge. Another difficulty is the implicit use of the concept of probability that does not appear anywhere during the quantum dynamics but only at the measurement instance, the exact time and the duration of which cannot be rigorously known.
@@
The application of the Born rule has become prevalent for the probabilistic interpretation of measurement outcomes. While this facilitates the use of classical concepts and language to describe quantum phenomena and properties, the trade-off is in the necessity to consider the measurement device as classical. Ceasing to apply the Born rule would allow to discard the notion of probabilities being a constitutive element of quantum mechanics thereby paving the way to its deterministic interpretation. The cost would then be a more elaborate description of the detection process. Given the widespread use of the theory in its probabilistic interpretation and its achievements, one may argue that while not minor, the matter remains a specialized problem with no strong consequence for experiments as discussed in a recent editorial [6]. However, Werner Heisenberg’s assertion that: “What we observe is not nature in itself but nature exposed to our method of questioning[7], can serve as a warning against the systematic use of any methodology that eludes fundamental difficulties to interpret experimental data.
@@ II The inadequacy of the Born rule for single events
At the time of the formulation of the Born rule in 1926 [3], physicists and chemists ambitioned and managed to gain a detailed knowledge and understanding of the electronic structures of atoms and molecules, neglecting the effects of the apparatus upon measurement of a system’s quantum properties. Contributions of other authors, including von Neumann[5], Dirac[8], and Jordan[9, 10], which adhere to the probabilistic interpretation of quantum mechanics, made the Born rule a postulate of quantum mechanics [11]. As this rule is still used nowadays, the dichotomy between the object of study and the measurement device persists; and it is legitimate to a why a detector during measurement is not described as a quantum system together with the object of study. Moreover, for single microscopic event (or individual) measurements, the use of this rule is questionable.
II.1 Redefinition of the measurement without the Born rule
@@
II.2 Deterministic view of quantum mechanics
@@
In this deterministic paradigm, the principle of realism amounts to saying that the measured macroscopic quantity has vanishing fluctuations. It sounds a priori idealistic that a perfect exact measurement requires an infinite number of interrogations but, by analogy, it sounds as idealistic to assume that space and time are continuous since one never approaches an infinitesimal distance or time increment. One needs only to accept that, in any modeling, concepts are defined only in some unreachable limit cases; otherwise, this would add another layer of unwanted complexity preventing a simple understanding of a physical phenomena. This corroborates the non-cloning theorem proposed by Zurek [1] on the impossibility to clone any unknown quantum state justifying the use of many copies to gain knowledge of this state before reproducing it. We therefore propose that: Quantum mechanics predicts deterministically that the humankind cannot predict everything.
@@
II.3 The reality of individual measurement
Ideally one seeks to optimally design experimental setups to mitigate measurement errors and minimize measurement uncertainty. This problem is typically resolved by carrying measurements over a large number of particles with beams of large number of photons to reduce the fluctuations over the measurement. But foundational experiments [14, 15, 16, 19, 17, 18, 20] studying individually one photon or an atom together with separate measurement events each identified as a “click” manifestation in a detector, suggest that a second revolution in quantum mechanics has taken place as one can infer from these experiments that quantum systems can be controlled at the individual level. One easily overlooked matter in this picture is that the so-called “click” signal corresponds in reality to a picture where an amplification of a single-particle signal has been transformed into a macroscopic outcome, as contrary to the latter the former cannot be detected by a human being. Another overlooked point is that, as shown further below, statistics over many events are always used in these “individual” experiments. Claims that each of these events is spatially and timely independent and quantum uncorrelated from each other contradict basic quantum features like uncertainty over position and momentum as well as over time and energy, and the indistinguishability of identical particles. Such claims need to be rigorously examined.
@@ III.2 Thermodynamics interpretation of the source
Since any particle beam source is always multi-mode and multi-particle, it is crucial to use a quantum field approach for a rigorous analysis of the detection process. Quite generally, the full quantum characterization of a source is impossible since it should be described within the framework of a Fock space by a global many-particle wave function $|\Psi\rangle$. For example, a source of gamma-ray is not point-like but a sample of matter made of a large number of radionuclides emitting gamma photons from wherever they are within the extent of the sample. If a spherical wave from the source can be expected on a statistical average, imperfections due to the finite spatial extension of the source generate a complex superposition of photon states that reveals by interference a localized surge of one or more particles that a detector can subsequently detect unambiguously (like in a bubble chamber or on a photographic plate for instance). Note that for high-energy particles, the uncertainty over the position is in principle very small and makes possible to treat the trajectory classically.
@@ III.3 Scattering approach
Quite generally, a photon detector can be modelled using scattering theory by analyzing the transition rate of excited electrons using the Fermi golden rule. This rule, which originates from the scattering theory using the S–matrix formalism, imposes an energy conservation between the initial and final states. Although these states look like non-localized states in space such as a pure monochromatic wave, the $i0^{+}$ prescription appearing in the Lippman-Schwinger equations guarantees an adiabatic switching on of a scattering process and imposes that any signal which is interacting with a detector should be causal, localized in time and in space, without any specific characterization of the extension of the wave function. In fact, a localized wave function always has an uncertainty in energy $\Delta E$, which implies to write it as a superposition of delocalized monochromatic input states, which are subsequently transformed into a localized superposition of delocalized output states under the action of S–matrix. As a result of this localisation, the uncertainty principle fixes a lower bound for the time of scattering given by the pulse time $t_{d}\sim\hbar/\Delta E$, which differs from the collision time related to the range of interaction.
@@ III.4 Determinism in photo-detector models
These models show the major limitations on the capabilities of detectors, which make it practically impossible with the current available technology to isolate at low energy one pulse event from another in order to control single photons well separated in time. The individual event approach in quantum mechanics works only in the statistical sense. As a result, the advantage of using an avalanche detector in comparison to a continuous detector easier to model, becomes less obvious. The above modelling changes totally the perception of the apparent paradoxes in quantum mechanics, in particular, Schrödinger’s cat and the Bell experiment.
@@ IV.2 The genuine single photon generation
Using a fainted coherent pulse seems a practical way of producing an isolated photon but justifying the absence of effects of higher photon numbers in the superposition of states, is an issue. By exciting a single “artificial” atom by means of a Rabi $\pi$-pulse from a classical radiation source, a subsequent spontaneous emission will produce a real single-photon pulse with a broadening given by the decay rate. Another way to produce one is to use entangled beams where the measurement of one photon in one beam “guarantees” the presence of one photon in another beam [19]. Owing to the precision of an avalanche diode; however, one may question the suitability of such a post-selection method once it is established that there cannot be separability between the events.
If this process is performed within a cavity, the single photon has a channel for producing a genuine superposition leading to a pure entangled state using a beam splitter: $|\psi\rangle_{1}=(c_{V}|0\rangle_{H}|1\rangle_{V}+c_{H}|1\rangle_{H}|0\rangle_{V})$ with $|c_{V}|^{2}+|c_{H}|^{2}=1$ as shown in Fig. 1. Quantum cloning is in principle possible in a known basis and therefore in the Fock basis. If we suppose that an ideal detector leads to the cat state from amplification: $|\psi\rangle_{n}=(c_{V}|0\rangle_{H}|n\rangle_{V}+c_{H}|n\rangle_{H}|0\rangle_{V})$, then the relative fluctuations are very large: $\sqrt{\langle\delta^{2}N_{\epsilon}\rangle}/{\langle\hat{N}_{\epsilon}\rangle}=1/\sqrt{|c_{\epsilon}|^{2}(1-|c_{\epsilon}|^{2})}$, and do not shrink for large cat sizes. But for a realistic case, such states do not appear to be created easily experimentally, since all fluorescence photons cannot be detected individually once we admit that the detection is not a markovian process.
@@
Figure 1: Example of non-Markovian detection based on model 2: Sequence of four $\pi/2$ pulses entering two avalanche detectors, resulting from four separate single photon signals passing through a 50/50 beam splitter (BS). Determinism imposes the absence of a quantum reset, which leads to a memory effect. Once the detector is in a $\pi$ state after the electronic transition, the avalanche is effective or not, according to the detector random noise state. If the two detectors are synchronized, the outcomes are correlated leading to coincidence measurement outcomes. If one detector state is shifted by $\pi/2$ relatively to the other which is more common on average, the outcomes are anti-correlated with no coincidence (see table) and corresponds to the perfect prediction of one photon per pulse if the avalanche is always effective. More generally, the impossibility of the detector to tell its own state due to its noise and the presence of many electrons (see model 3), instead, prevent any synchronisation causing the randomness of both single and coincidence outcomes such as in Ref. [19]. In contrast, a model detector based on Born’s rule would have produced 0/1 or 1/0 outcomes with probability 1/2.

### v2

@@ Abstract
We argue that measurement data in quantum physics can be rigorously interpreted only as a result of a statistical, macroscopic process, taking into account the indistinguishable character of identical particles. Quantum determinism is in principle possible on the condition that a fully-fledged quantum-field-theoretic model is used to describe the measurement device in interaction with the studied object as one system. In contrast, any approach that relies on Born’s rule discriminates the dynamics of a quantum system from that of the detector with which it interacts during measurement. In this work, we critically analyze the validity of this measurement postulate applied to single-event signals. In fact, the concept of “individual” particle becomes inadequate once both indistinguishability and a scattering approach allowing an unlimited interaction time for an effective detection, are considered as they should be, hence preventing the separability of two successive measurement events. In this context, measurement data should therefore be understood only as a result of statistics over many events. Accounting for the intrinsic noise of the sources and the detectors, we also show with the illustrative cases of the Schrödinger cat and the Bell experiment that once the Born rule is abandoned on the level of a single particle, realism, locality and causality are restored. We conclude that indiscernibility and long-time detection process make quantum physics not fundamentally probabilistic.
@@ I Introduction
Here, we argue that the Born rule (or probability measurement postulate) is a non-essential mathematical convenience that should no longer be used, even though for nearly 100 years, much research in quantum physics has relied upon it for the interpretation of quantum experimental data. In a nutshell, this rule states that the probability density of finding a system in a given state, immediately after measurement, is proportional to the square of the amplitude of the system’s wave function in that state immediately before measurement, thus assuming that the system was isolated from its environment during the measurement process. In this framework, the measurement of a physical quantity can only yield one of the eigenvalues of the corresponding observable with a certain probability. As such, the Born rule separates the dynamics of the studied quantum system from that of the measurement device, thus assuming that measurement data correspond to observables as properties of the studied quantum system only. However, the elephant in the room that has been systematically overlooked in practice, is that there is no reason why quantum mechanics should discriminate between the object of study and the measurement device it is in interaction with as though they were independent entities. A precise knowledge of the interaction between the detector and the quantum system is required for a detailed and rigorous description of the measurement process and hence a sound interpretation of the data. This is a considerable challenge.
The separation between the quantum or microscopic world and the macroscopic one that includes measurement devices stems from quantum indeterminacy along with the so-called “Copenhagen interpretation” (regardless of its many “flavors” as Mermin put it [6]). More precisely, in this interpretation, separation appears to be necessary because while the wave function of an isolated of quantum system can be defined as a linear superposition of eigenfunctions that satisfy the Schrödinger equation, there is no such linear combination that may apply for macroscopic systems (measurement device) states as the classical equations of motion such as, e.g., the Hamilton-Jacobi or Lagrange, are nonlinear. Hence the pervasive concept of probability in the description of quantum phenomena as indeterministic in nature, while probabilities do not appear anywhere during the quantum dynamics but only at the measurement instance, the exact time and the duration of which cannot be rigorously known. Consequently, in this paradigm, an observation or measurement of a quantum system’s property is an irreversible stochastic process that leads to the collapse of the wave function into one of the isolated (before measurement) system’s eigenstates. The resulting wave function after measurement is interpreted as a quantum back action [7] and is subsequently governed by the unitary quantum dynamics.
It is important to point out here that, as noted in Ref. [8], there is no certain knowledge about the destruction or preservation of the other eigenstates of the state space after the collapse. None of the foundational works of Born [3], Dirac [4], von Neumann [5] and, later, Lüders [9] provide a formal proof of the wave function collapse, so one should simply accept that, as already pointed above, properties of quantum systems only “secure” their actual value upon measurement, and by that means become physical quantities. Several notable works however, have questioned the Copenhagen interpretation and the wave function collapse, and have provided their own explanation. These include the de Broglie-Bohm theory [10, 11, 12], Everett’s “many-worlds” interpretation [13], the transactional interpretation [14], ’t Hooft’s cellular automaton interpretation [15], and several others discussed in, e.g., Refs. [8, 16]. Here, for brevity, we simply note that the de Broglie–Bohm theory, which is non-local yet deterministic, adds a layer of complexity by introducing the pilot equation and becomes difficult to extend in a relativistic formulation. And, while the many-worlds interpretation does without Born’s rule, it “splits” the universe into mutually unobservable alternate histories, which can hardly be proven experimentally.
@@
As stated in Ref. [16], either of the following three courses of action can in principle be chosen to manage the conceptual difficulties that plague the interpretation of actual measurement data: i/ a change of interpretation of quantum mechanics; ii/ a change of its mathematical formulation; iii/ a modification of the theory. As of yet, no breakthrough is in sight, and the application of the Born rule that fostered the probabilistic interpretation of measurement outcomes remains prevalent. While this facilitates the use of classical concepts and language to describe quantum phenomena and properties, the trade-off is in the necessity to consider the measurement device as classical [15]. Ceasing to apply the Born rule would allow to discard the notion of probabilities being a constitutive element of quantum mechanics thereby paving the way to its deterministic interpretation. The cost would then be a more elaborate description of the detection process. Given the widespread use of the theory in its probabilistic interpretation and its achievements, one may argue that while not minor, the matter remains a specialized problem with no strong consequence for experiments as discussed in a recent editorial [17]. However, Werner Heisenberg’s assertion that: “What we observe is not nature in itself but nature exposed to our method of questioning[18], can serve as a warning against the systematic use of any methodology that eludes fundamental difficulties to interpret experimental data.
@@ II The inadequacy of the Born rule for single events
II.1 Obstacles on the path towards the second revolution
At the time of the formulation of the Born rule in 1926 [3], physicists and chemists ambitioned and managed to gain a detailed knowledge and understanding of the electronic structures of atoms and molecules, neglecting the effects of the apparatus upon measurement of a system’s quantum properties. Contributions of other authors, including von Neumann [5], Dirac [19], and Jordan [20, 21], which adhere to the probabilistic interpretation of quantum mechanics, made the Born rule a postulate of quantum mechanics [22]. These works were key to move from the early or “old” quantum mechanics to its modern formulation; but to foster the development of quantum technology, a second revolution of quantum mechanics [25, 26] has taken place with new theoretical and experimental methods to design and manufacture the next generation of useful quantum devices and processes for, e.g., quantum computing, quantum cryptography, quantum teleportation, and precision metrology. These methods aim to allow manipulation and control of single atoms, photons and ions [27, 28]. As far as measurements are concerned, a huge host of modern literature is devoted to the field, focusing in particular on quantum trajectories, continuous measurements, weak or generalized (ancilla-based) measurements, weak values, and the effect on environment on measurements in open systems [29, 30, 31, 32, 33, 34, 35, 36, 37]. We highlight here that none of these representative works related to measurement are based on a quantum field theory, thus ignoring the formalism of the Fock space for the interpretation of measurement outcomes in the quantum state space fully accounting for quantum statistics and long-time detection.
Further, the joint system-detector evolution and the corresponding back-action effects are nowadays one of the cornerstones of the measurement-based quantum state engineering, performed experimentally on various quantum-hardware platforms [7]. Even so, these developments are based on rules akin to the Born rule (Lüders’ and von Neumann’s projection postulates, and stochastic approach) [38, 39], lacking the heuristic approach necessary to grasp the whole physics aspects associated to a physical phenomenon. It misses the fundamental question as to why quantum randomness should be considered as an irreducible feature of quantum physics. Without a measurement procedure necessary for quantum engineering and made by humans, any quantum unitary evolution of an isolated system does not made any use of randomness. It is therefore the simplifying assumptions (trace of the noise bath and detector description) that renders unavoidable the use of probabilities.
In light of the above, the two assertions related to randomness in Ref. [25]:
Uncertainty principle: for every perfectly specified quantum state there is always at least one measurement, the results of which are completely certain, and simultaneously at least one measurement for which the results are largely random.
and
Decoherence: what happens to quantum superpositions when an attempt is made to distinguish previously indistinguishable ways an event can be realized. It renders superpositions of probability amplitudes into superpositions of classical probabilities. Decoherence has no analogue in classical physics.
are therefore questionable as there is no proof that randomness is an irreducible characteristics of quantum physics. In the first assertion, the quantum uncertainty associated to the position and momentum of a particle should not be considered as evidence of the probabilistic nature of quantum mechanics. Position and momentum are concepts that belong to classical physics but loose relevance and rigor at the quantum scale, especially on the level of single particles. In the second assertion, the reason for randomness is in the simplifying assumption that consists in adding a thermal or quantum noise in the modelling (e.g., some leakage in a source or in a beam, or during the detection). Formally, it involves a projection technique analogue to that used when applying Born’s rule. These assertions support non-locality as an essential feature of quantum mechanics and hence lead to incongruities such as, e.g., the EPR paradox [40], the quantum eraser paradox [41] or Wheeler’s delayed-choice experiment [42], since all the literature associated to the second revolution assumes the reality of a single detection event without accounting for indistinguishability and long time detection.
With the Born rule remaining at the heart of most theoretical approaches in quantum physics, the dichotomy between the object of study and the measurement device persists; and it is legitimate to ask why a detector during measurement is not described as a quantum system together with the object of study. Moreover, for single microscopic event (or individual) measurements, the use of this rule is questionable.
II.2 Redefinition of the measurement without the Born rule
@@
II.3 Deterministic view of quantum mechanics
@@
In this deterministic paradigm, the principle of realism amounts to saying that the measured macroscopic quantity has vanishing fluctuations. It sounds a priori idealistic that a perfect exact measurement requires an infinite number of interrogations but, by analogy, it sounds as idealistic to assume that space and time are continuous since one never approaches an infinitesimal distance or time increment. One needs only to accept that, in any modeling, concepts are defined only in some unreachable limit Leggettcases; otherwise, this would add another layer of unwanted complexity preventing a simple understanding of a physical phenomena. This corroborates the non-cloning theorem proposed by Zurek [1] on the impossibility to clone any unknown quantum state justifying the use of many copies to gain knowledge of this state before reproducing it. We therefore propose that: Quantum mechanics predicts deterministically that the humankind cannot predict everything.
@@
II.4 The reality of individual measurement
Ideally one seeks to optimally design experimental setups to mitigate risks of measurement errors and minimize measurement uncertainty. This problem is typically resolved by carrying measurements over a large number of particles with beams of large number of photons to reduce the fluctuations over the measurement. But foundational experiments [45, 46, 47, 50, 48, 49, 51] studying individually one photon or an atom together with separate measurement events each identified as a “click” manifestation in a detector, suggest that a second revolution in quantum mechanics has taken place as one can infer from these experiments that quantum systems can be controlled at the individual level. One easily overlooked matter in this picture is that the so-called “click” signal corresponds in reality to a picture where an amplification of a single-particle signal has been transformed into a macroscopic outcome, as contrary to the latter the former cannot be detected by a human being. Another overlooked point is that, as shown further below, statistics over many events are always used in these “individual” experiments. Claims that each of these events is spatially and timely independent and quantum uncorrelated from each other contradict basic quantum features like uncertainty between position and momentum, as well as over time and energy, and the indistinguishability of identical particles. Such claims need to be rigorously examined.
@@ III.2 Thermodynamics interpretation of the source
Since any particle beam source is always multi-mode and multi-particle, it is crucial to use a quantum field approach for a rigorous analysis of the detection process. Quite generally, the full quantum characterization of a source is impossible since it should be described within the framework of a Fock space by a global many-particle wave function $|\Psi\rangle$. For example, a source of gamma-ray is not point-like but a sample of matter made of a large number of radionuclides emitting gamma photons from wherever they are within the extent of the sample. If a spherical wave from the source can be expected on a statistical average, imperfections due to the finite spatial extension of the source generate a complex superposition of photon states that reveals by interference a localized surge of one or more particles that a detector can subsequently detect unambiguously (like in a bubble chamber or on a photographic plate for instance). Note that for high-energy particles, the uncertainty over the position is in principle very small and makes possible to treat the trajectory classically [52].
@@ III.3 Scattering approach
Quite generally, a photon detector can be modelled using scattering theory by analyzing the transition rate of excited electrons using the Fermi golden rule. This rule, which originates from the scattering theory using the S–matrix formalism, imposes an energy conservation between the initial and final states. Although these states look like non-localized states in space such as a pure monochromatic wave, the $i0^{+}$ prescription appearing in the Lippman-Schwinger equations guarantees an adiabatic switching on of a scattering process and imposes that any signal which is interacting with a detector should be causal, localized in time and in space, without any specific characterization of the extension of the wave function. In fact, a localized wave function always has an uncertainty in energy $\Delta E$, which implies to write it as a superposition of delocalized monochromatic input states, which are subsequently transformed into a localized superposition of delocalized output states under the action of S-matrix. As a result of this localisation, the uncertainty principle fixes a lower bound for the time of scattering given by the pulse time $t_{d}\sim\hbar/\Delta E$, which differs from the collision time related to the range of interaction.
@@ III.4 Determinism in photo-detector models
These models show the major limitations of the detectors capabilities, which make it practically impossible with the current available technology to isolate at low energy one pulse event from another in order to control single photons well-separated in time. Moreover, one common bias in the analysis of experiments is to expect that if the statistical distribution of particles in a beam is Maxwellian, the particles behave independently and classically while even in this case, indistinguishability still plays a role during detection. However, indistinguishability cannot be directly observed as it involves the detector itself, and one would need a second detector measuring the state of the first one before it measures any signal (but to our knowledge, such a realization has not yet been implemented). Hence, the individual event approach in quantum mechanics works only in the statistical sense. As a result, the advantage of using an avalanche detector in comparison to a continuous detector easier to model, becomes less obvious. The above modeling approach changes totally the perception of the apparent paradoxes in quantum mechanics, in particular, Schrödinger’s cat and the Bell experiment, discussed below.
@@ IV.2 The genuine single photon generation
Using a fainted coherent pulse seems a practical way of producing an isolated photon but justifying the absence of effects of higher photon numbers in the superposition of states, is an issue. By exciting a single “artificial” atom by means of a Rabi $\pi$-pulse from a classical radiation source, a subsequent spontaneous emission will produce a real single-photon pulse with a broadening given by the decay rate. Another way to produce one is to use entangled beams where the measurement of one photon in one beam “guarantees” the presence of one photon in another beam [50]. Owing to the precision of an avalanche diode; however, one may question the suitability of such a post-selection method once it is established that there cannot be separability between the events. If this process is performed within a cavity, the single photon has a channel for producing a genuine superposition leading to a pure entangled state using a beam splitter: $|\psi\rangle_{1}=(c_{V}|0\rangle_{H}|1\rangle_{V}+c_{H}|1\rangle_{H}|0\rangle_{V})$ with $|c_{V}|^{2}+|c_{H}|^{2}=1$ as shown in Fig. 1. Quantum cloning is in principle possible in a known basis and therefore in the Fock basis. If we suppose that an ideal detector leads to the cat state from amplification: $|\psi\rangle_{n}=(c_{V}|0\rangle_{H}|n\rangle_{V}+c_{H}|n\rangle_{H}|0\rangle_{V})$, then the relative fluctuations are very large: $\sqrt{\langle\delta^{2}N_{\epsilon}\rangle}/{\langle\hat{N}_{\epsilon}\rangle}=1/\sqrt{|c_{\epsilon}|^{2}(1-|c_{\epsilon}|^{2})}$, and do not shrink for large cat sizes. But for a realistic case, such states do not appear to be created easily experimentally since all fluorescence photons cannot be detected individually; once we admit that the detection is not a markovian process.
@@
Figure 1: Example of non-Markovian detection based on model 2: Sequence of four $\pi/2$ pulses entering two avalanche detectors, resulting from four separate single photon signals passing through a 50/50 beam splitter (BS). Determinism imposes the absence of a quantum reset, which leads to a memory effect. Once the detector is in a $\pi$ state after the electronic transition, the avalanche is effective or not, according to the detector random noise state. If the two detectors are synchronized, the outcomes are correlated leading to coincidence measurement outcomes. If one detector state is shifted by $\pi/2$ relatively to the other which is more common on average, the outcomes are anti-correlated with no coincidence (see table) and corresponds to the perfect prediction of one photon per pulse if the avalanche is always effective. More generally, the impossibility of the detector to tell its own state due to its noise and the presence of many electrons (see model 3), instead, prevent any synchronisation causing the randomness of both single and coincidence outcomes such as in Ref. [50]. In contrast, a model detector based on Born’s rule would have produced 0/1 or 1/0 outcomes with probability 1/2. In this example, the decoherence time for the detector to reset or not to its ground state between two events (like two separated pulses of light) is one factor responsible for this randomness.